The area of a rectangle is given by A = x2 – 2x + 30. Find the value of x if the total area of the rectangle is equal to 45 square units.
Answer by best82394(4)

If the area of the rectangle is 45 units, then x^2-2x+30=45. Subtract 30: x^2-2x=15. Bring the 15 to the other side, and then factor. (x-5)(x+3)=0 Then you must CHECK your answer, and you will see that (x+3) does not work. So (x-5) is the only one that works. (x-5)=0, zero product theorem. therefore, the value of x that satisfies the problem is 5.